8. You deposit $1000 in an account today You will deposit $600 at the end of each month for the next twelve months and $800 per month for the following twelve months. How much interest will you have earned in two years in the account pays 5.5% compounded monthly?
Answer #1
EAR = [(1 +stated rate/no. of compounding periods) ^no. of compounding periods - 1]* 100
? = ((1+5.5/(12*100))^12-1)*100
Effective Annual Rate% = 5.6408
Future value = present value*(1+ rate)^time
Future value = 1000*(1+0.056408)^2
Future value = 1116
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 600*(((1+ 5.5/1200)^(1*12)-1)/(5.5/1200))
FV = 7384.3
Future value = present value*(1+ rate)^time
Future value = 7384.3*(1+0.056408)^1
Future value = 7800.83
FVOrdinary Annuity = C*(((1 + i )^n -1)/i)
C = Cash flow per period
i = interest rate
n = number of payments
FV= 800*(((1+ 5.5/1200)^(1*12)-1)/(5.5/1200))
FV = 9845.74

total = 9845.74+7800.83+1116

=

18762.57